Swifty Appliance Center is an experienced home appliance dealer. Swifty Appliance Center also offers a number of services for the home appliances that it sells. Assume that Swifty Appliance Center sells ovens on a standalone basis. Swifty Appliance Center also sells installation services and maintenance services for ovens. However, Swifty Appliance Center does not offer installation or maintenance services to customers who buy ovens from other vendors. Pricing for ovens is as follows.

Oven only $800
Oven with installation service 870
Oven with maintenance services 960
Oven with installation and maintenance services 990


In each instance in which maintenance services are provided, the maintenance service is separately priced within the arrangement at $160. Additionally, the incremental amount charged by Swifty Appliance Center for installation approximates the amount charged by independent third parties. Ovens are sold subject to a general right of return. If a customer purchases an oven with installation and/or maintenance services, in the event Swifty Appliance Center does not complete the service satisfactorily, the customer is only entitled to a refund of the portion of the fee that exceeds $800.

Assume that a customer purchases an oven with both installation and maintenance services for $990.

(b) Indicate the amount of revenue that should be allocated to the oven, the installation, and to the maintenance contract. (Do not round intermediate calculations. Round final answers to 2 decimal places.)

Oven $

Installation $

Maintenance $
Answer #1

ANSWER:

Allocation
oven    800 800*990/1030= 768.93
Installation 70 70*990/1030= 67.28
Maintenance 160 160*990/1030= 153.78
1030 990

when oven is sold with maintenance then price of oven = 960-160 = 800

oven when sold with installation : 870-800 =70 is the cost of installation

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************